# Descriptive Image Quality Assessment in the Wild

### v2

@@ 5.2 Results on Benchmarks
Quantitative results of assessment reasoning and comparison reasoning are illustrated in Tab. 5. First, the performance of the two VLM-specific models significantly declines when applied to tasks outside their defined scopes. For instance, Q-Instruct is not good at comparison tasks and Co-Instruct’s performance is unsatisfactory on full-reference tasks. Second, GPT-4V shows robust reasoning abilities, stably outperforming prior IQA-specific VLMs. Third, DepictQA-Wild still surpasses GPT-4V, especially in the non-reference setting, affirming its superior reasoning capabilities.
@@
Table 5: Assessment reasoning and comparison reasoning results under both single-distortion and multi-distortion cases. GPT-4 score metric is reported in the full-reference / non-reference setting.
@@
GPT-4V [41] 67.8 / 59.2 71.0 / 62.3 66.2 / 60.3 67.2 / 60.1
Q-Instruct [62] 46.8 / 45.7 48.2 / 45.8 35.3 / 35.6 33.7 / 36.0
Co-Instruct [64] 40.1 / 45.3 41.9 / 46.7 37.6 / 48.1 35.6 / 48.0
DepictQA-Wild 76.8 / 74.2 75.6 / 72.3 75.1 / 74.9 71.7 / 68.7
Figure 4: User study.
@@
Qualitative results of our model on the four tasks in the non-reference setting are depicted in Fig. 2. More qualitative results are provided in Appendix.
5.3 Real-world Applications
Assessing web-downloaded images. A practical usage of an IQA model involves assessing the quality of real images. We collect a total of 50 real-world images from the web, featuring diverse contents including animals, plants, faces, buildings, and landscapes. Qualitative results in Fig. 5 indicate that our method can assess real images with detailed descriptions. More importantly, DepictQA-Wild can describe how the distortions affect the contents. For example, in Fig. 5d, our model first accurately identifies the “severe quantization”, then describes that the quantization “causes banding in the sky and water”, and finally concludes that the quality “is considerably degraded”. We also conduct a user study with 20 participants involved. Participants are instructed to choose the assessment result that is of the highest quality among the test methods. The results are shown in Fig. 4, revealing that our approach stably outperforms baseline methods in aligning human perception.
Figure 5: Qualitative results on assessing web-downloaded images.
Table 6: Results on model-processed images.
NIQE ClipIQA MUSIQ ManIQA GPT-4V DepictQA-Wild
Rank $\downarrow$ 2.20 1.40 1.60 1.80 1.34$\pm$0.27 1.20
Accuracy $\uparrow$ 45.5 72.7 77.3 66.4 74.5 82.7
Comparison on model-processed images. To develop image restoration models, one often needs to compare the restoration quality of different models. To simulate this, we consider five distortions including “defocus blur”, “motion blur”, “noise”, “JPEG compression”, and “low resolution”. For each distortion, three to four candidate models are used to process the distorted images. We manually rank the restored results, assigning “1” to the best restoration, “2” to the second best, etc. Different IQA methods are adopted to compare these restored images pairwise and find the best restoration. The average rank of the found best restoration and the accuracy of the paired comparison are reported in Tab. 6. First, DepictQA-Wild achieves an average rank of 1.20 (1 is the best), outperforming both GPT-4V and score-based methods. Second, though the temperature is set to 0, GPT-4V shows variability with a large standard deviation. Third, model-restored images are generally out-of-distribution for our model, while DepictQA-Wild exhibits excellent generalization ability on these images.
@@ 5.3 Real-world Applications
Table 7: Our assistant model surpasses GPT-4V greatly in instant rating task. The metric is accuracy in the full-reference / non-reference setting.
GPT-4V Our Assistant
TID2013 84.4 / 88.1 94.9 / 94.6
LIVE-MD 79.6 / 72.7 93.1 / 92.8
MDID2013 70.6 / 67.6 90.1 / 89.8
@@
Table 8: Retaining resolution is important to identify the images with better aspect ratio or higher resolution.
@@
73.0 91.7 77.2
85.6 99.0 94.8
98.8 99.3 96.8
Figure 6: Confidence and performance are highly correlated.
5.4 Ablation Studies
Assistant model. To construct comparison reasoning responses, we train an assistant model to predict comparison results (see Fig. 3). These results serve as pseudo labels, which are subsequently provided to GPT-4V to generate responses. We compare the assistant model to GPT-4V on three out-of-distribution IQA datasets. The results in Tab. 7 affirm the superiority of the assistant model.
@@ 5.4 Ablation Studies
Confidence. We examine the correlation between model performance and estimated confidence in Fig. 6. For distortion identification and instant rating tasks, across both full-reference and non-reference settings, our model demonstrates improved performance as the confidence interval increases. These findings validate the effectiveness of our confidence estimation. More results in Appendix.
6 Conclusions and Limitations
@@
Limitations. First, the fine-grained abilities, which require more high-level perception skills, are still not satisfactory. For example, in Fig. 5c, though identifying noise and pixelation successfully, our model fails to point out that they are respectively located in the left and right parts. Second, the task paradigm is not fully unified. Our comparison tasks focus on images with the same content but different distortions. Extending this to include comparisons among different contents could enhance the model’s applicability. Furthermore, besides quality, image aesthetics can also be described linguistically based on VLMs. Third, whether our assessment can be used as feedback to improve the quality of generation or restoration models is still under-explored. These belong to our future works.
Appendix
Appendix A Overview
This Appendix is structured as follows. Dataset details are described in Appendix B, followed by the details of model, training and inference in Appendix C. More ablation studies, qualitative results, and the details of real-world applications are presented in Appendix D.
Appendix B Dataset Details
B.1 Details of Distortion Library
Figure A1: Reference image for all distortion examples shown in Fig. A2.
@@ B.1 Details of Distortion Library
As stated in Sec. 3.2, to facilitate the dataset construction, we design and implement a comprehensive distortion library. Our distortion system contains 12 distortion super-categories in total, with each category consisting of multiple sub-categories. For instance, the “blur” category encompasses “Gaussian blur”, “motion blur”, “lens blur”, etc. In total, there are 35 sub-categories. For each sub-category, there are 5 severity levels: “slight”, “moderate”, “obvious”, “serious”, and “catastrophic”. In this section, we elaborate on our distortion implementations [44], including the principles, formulas, and severity setup.We also provide one example for each implementation in Fig. A2, with the reference image in Fig. A1.
@@
  • Glass blur. Filter the image using a Gaussian filter, then randomly jitter each pixel in the image by $x$ pixels, and repeat this process $n$ iterations. $[\sigma,x,n]\in[(0.7,1,1),(0.9,2,1),(1.2,2,2),(1.4,3,2),(1.6,4,2)]$.
@@
  • Gaussian noise in YCrCb space. Similar to the Gaussian noise in RGB space, this distortion is implemented in YCbCr space, where $(\sigma_{l},\sigma_{r},\sigma_{b})\in[(0.05,1,1),(0.06,1.45,1.45),(0.07,1.9,1.9),(0.08,2.35,2.35),(0.09,2.8,2.8)]$.
@@
[…]
@@
Figure A2: Distortion examples of our distortion design. We showcase one example for each distortion implementation. The reference image is depicted in Fig. A1.
Multi-distortion setups. As discussed in Sec. 3.2, multiple distortions may occur simultaneously on the same image in practical usage. First, we observe that humans can identify at most two distortions when three or more are applied, thus we limit the number of applied distortions to two. Second, some distortions could weaken each other’s presentation (e.g., “brighten” weakens “darken”, “blur” weakens “over-sharpen”). Also, certain distortions show similar visual effects (e.g., “pixelate” looks similar to “blur”), making it hard to identify both if applied simultaneously. Hence, to exclude contradictory or similar distortion combinations, we manually review all possible combinations, and the results in Tab. A1 present all feasible distortion combinations used in our dataset construction.
Table A1: Multi-distortion setting where we show all feasible distortion combinations.
@@
B.2 Details of Dataset Construction

### v3

@@ V-B Results on Benchmarks
Quantitative results of assessment reasoning and comparison reasoning are shown in Tab. V and Tab. VI. First, the performance of the VLM-specific models significantly declines on tasks outside their defined scopes. For instance, Co-Instruct’s performance is unsatisfactory on full-reference tasks. Second, GPT-4V shows robust reasoning abilities, stably outperforming prior IQA-specific VLMs. Third, DepictQA-Wild surpasses GPT-4V, especially in the non-reference setting, affirming its superior reasoning abilities. Finally, DepictQA-Wild achieves relatively good GPT-4 score and ROUGE-L, indicative of the overall semantic accuracy, but a low BLEU score (yet remains much higher than GPT-4V), which reflects word-level consistency. This suggests that while our predicted answers do not precisely duplicate the ground truths word-for-word, they preserve similar meanings with diverse expressions.
@@
TABLE V: Assessment reasoning and comparison reasoning results with GPT-4 score metric in the full-reference / non-reference setting. The best results are marked in bold.
Methods Assessment Reasoning Comparison Reasoning
Single-dist. Multi-dist. Single-dist. Multi-dist.
GPT-4V 67.8 / 59.2 71.0 / 62.3 66.2 / 60.3 67.2 / 60.1
Q-Instruct - / 45.7 - / 45.8 - -
Co-Instruct 40.1 / 45.3 41.9 / 46.7 37.6 / 48.1 35.6 / 48.0
DepictQA-Wild 76.8 / 74.2 75.6 / 72.3 75.1 / 74.9 71.7 / 68.7
TABLE VI: Assessment reasoning and comparison reasoning results with classical metrics (BLEU and ROUGE-L) in the full-reference / non-reference setting. The best results are bold.
@@
BLEU ROUGE-L BLEU ROUGE-L BLEU ROUGE-L BLEU ROUGE-L
GPT-4V 0.020/0.010 0.248/0.224 0.023/0.015 0.246/0.223 0.029/0.025 0.261/0.243 0.030/0.024 0.251/0.238
Q-Instruct - / 0.003 - / 0.210 - / 0.002 - / 0.198 - - - -
Co-Instruct 0.008/0.005 0.201/0.204 0.002/0.003 0.209/0.203 0.039/0.041 0.239/0.234 0.034/0.036 0.239/0.234
Ours 0.132/0.129 0.423/0.422 0.180/0.170 0.420/0.415 0.207/0.207 0.466/0.463 0.176/0.172 0.420/0.413
Figure 7: User study results.
@@
Qualitative results of our model on the four tasks in the non-reference setting are depicted in Fig. 2. More qualitative results are provided in Figs. S4, S5, S6, S7 and S8 of Supp. Mat.
V-C Quality score regression
Our key focus in this work is to generate descriptive language rather than quality scores. We focus more on linguistic descriptions because language is an effective interaction tool in an LLM-based intelligent agent. With the rapid development of LLMs and multi-modal techniques, in an LLM-based intelligent agent, language could be a useful tool for interacting and communicating across quality-related tasks such as image assessment, refinement, editing, and recommendation. Still, if it is required, our approach can produce quality scores.
Quality score regression. The score regression results are evaluated on the PIPAL, KADID, TID2013, and CSIQ datasets. These datasets include high-quality reference images and their distorted versions under various distortions. We calculate the win rate of an image against others to determine its quality score. Specifically, for an image A, we randomly sample comparison candidates, such as B, C, D, etc., which share the same content as A but have different distortions. Image A is then compared pairwise with each of its comparison candidates (B, C, D, etc.). Finally, the win rate of Image A against all its compared candidates is calculated as its quality score. The comparison numbers per image for PIPAL, KADID, TID2013, and CSIQ datasets are 58, 62, 60, and 15, respectively. We show that the comparison number per image could be reduced significantly without large performance degradation in Tab. S3 of Supp. Mat. The results of quality score regression are given in Tab. VII and Tab. VII, proving that our method can generate accurate quality scores.
TABLE VII: Results of quality score regression with SRCC / PLCC metrics. The best results are emphasized in bold.
[…]
Assessing in-the-wild images with different contents. Existing real-world IQA datasets like KonIQ [koniq] and SPAQ [spaq] contain real distorted images with various contents. To regress quality scores from such a dataset, our model needs to compare images with different contents though it is trained only to compare images with similar contents, as shown in Task 2 of Fig. 2. The results in Tab. VIII show that even with a task gap between training and test, our original DepictQA-Wild still achieves comparable results with previous score-based IQA methods in score regression. Furthermore, we formulate real-world IQA datasets into instant rating tasks to re-train our DepictQA-Wild, i.e., trained on KonIQ then evaluated on SPAQ, and vice versa. Our re-trained DepictQA-Wild outperforms all baselines trained on the same dataset. These results indicate that our method is capable of assessing in-the-wild images with different contents.
TABLE VIII: Results of quality score calculation on SPAQ and KonIQ datasets with SRCC / PLCC metrics. The best results are bold. DepictQA-Wild needs to compare images with different contents to obtain the quality score, since all images in the two datasets contain different contents. The original DepictQA-Wild is only trained to compare images with similar contents, which brings a task gap. When trained on the same dataset as baselines, DepictQA-Wild surpasses the baseline methods.
(a) Results on SPAQ dataset. Methods NIQE CLIPIQA MUSIQ MANIQA Ours Ours Train Set - - KonIQ KonIQ KonIQ Original SRCC 0.664 0.700 0.856 0.755 0.859 0.835 PLCC 0.679 0.722 0.859 0.765 0.861 0.841
(b) Results on KonIQ dataset. Methods NIQE CLIPIQA DBCNN MUSIQ Ours Ours Train Set - - SPAQ SPAQ SPAQ Original SRCC 0.530 0.685 0.731 0.753 0.787 0.717 PLCC 0.533 0.717 0.758 0.680 0.807 0.729
@@ V-C Quality score regression
TABLE IX: Our assistant model surpasses GPT-4V in instant rating task. The metric is accuracy in the full-reference / non-reference setting. The best results are highlighted in bold.
TID2013 LIVE-MD MDID2013
GPT-4V 84.4 / 88.1 79.6 / 72.7 70.6 / 67.6
Our Assistant Model 94.9 / 94.6 93.1 / 92.8 90.1 / 89.8
Figure 8: One example of model-restored image by SwinIR [SwinIR].
TABLE X: Relationships between the comparison reasoning and instant rating tasks. The best results are bold. Overall, the two tasks are beneficial to each other.
(a) Results on the instant rating task. BAPPStest KADIDtest PIPALtest TID2013 LIVE-MD MDID2013 Only Rating 81.6 / 81.6 92.4 / 92.3 89.1 / 89.0 94.2 / 94.1 92.9 / 92.7 92.1 / 91.7 Co-training 84.7 / 82.4 93.6 / 93.1 90.5 / 90.0 96.9 / 96.4 92.1 / 91.8 90.0 / 89.6
(b) Results on the comparison reasoning task. Single-distortion Multi-distortion GPT-4 Score BLEU ROUGE-L GPT-4 Score BLEU ROUGE-L Only Reasoning 74.3 / 69.6 0.203 / 0.202 0.465 / 0.453 70.6 / 69.1 0.165 / 0.165 0.414 / 0.407 Co-training 75.1 / 74.9 0.207 / 0.207 0.466 / 0.463 71.7 / 68.7 0.176 / 0.172 0.420 / 0.413
Results on Q-Bench. Q-Bench [qbench] is a VLM benchmark for low-level perception, but it mainly contains multiple-choice questions, whereas our model targets descriptive assessments, making direct evaluation mismatched. To test whether our data still helps on Q-Bench, we co-train with the multiple-choice dataset Q-Instruct [qinstruct] in Tab. XI. Co-training on Q-Instruct and our dataset consistently outperforms training on Q-Instruct alone, especially in distortion-related questions, demonstrating the benefit of our dataset even under a format shift. We borrow LLaVA results trained on Q-Instruct from [qinstruct] for reference.
TABLE XI: Results on Q-Bench [qbench]. The best results are highlighted in bold.
Q-Bench Yes-or-No What How Distortion Other
LLaVA (Q-Instruct trained) 66.4 58.2 50.5 49.4 65.7
Ours (Q-Instruct trained) 70.5 49.6 51.7 55.4 56.7
Ours (Q-Instruct + Our data trained) 74.0 62.8 54.0 66.7 63.9
V-D Ablation Studies
Relationships between the comparison reasoning and instant rating tasks are studied in Tab. X. First, comparison reasoning task improves the performance on four instant rating datasets, but decreases the results on two datasets. Overall, comparison reasoning task helps the instant rating. Second, instant rating task stably improves the performance on comparison reasoning task.
Figure 9: Our estimated confidence scores are high correlated to the model performance on (a) distortion identification and (b) (c) instant rating tasks on different benchmarks in both full-reference and non-reference settings.
Confidence. We examine the correlation between the performance and estimated confidence in Fig. 9. For distortion identification and instant rating tasks, across both full-reference and non-reference settings, our model demonstrates improved performance as the confidence interval increases. This validates the effectiveness of our confidence estimation.
Assistant model. To construct comparison reasoning responses, we train an assistant model to predict comparison results (see Fig. 3). These results serve as pseudo labels, which are subsequently provided to GPT-4V to generate responses. We compare the assistant model to GPT-4V on three out-of-distribution IQA datasets. The results in Tab. IX affirm the superiority of the assistant model.
[…]
@@ V-D Ablation Studies
TABLE XII: Retaining resolution during both training and inference is important to identify images with better aspect ratio or higher resolution.
@@
73.0 99.0 93.5 91.7 83.8 77.2 71.2
85.6 99.8 99.4 99.0 95.9 94.8 89.4
98.8 99.9 99.6 99.3 99.1 96.8 97.0
Ablation study on module choice. Under the same VLM architecture framework and the same training loss, we compare two vision-text connectors (i.e., vision abstractor v.s. projector) and three LLMs on distortion identification and instant rating tasks. The default vision-text connector and LLM in this ablation study are vision projector and Vicuna-v1.5-7B. The results in Tab. XIII show that simply replacing vision-text connectors or LLMs has relatively little influence on performance. Considering that vision abstractor can greatly reduce the computational burden than projector, we select abstractor by default. For example, for a $448\times 448$ image, projector generates 1024 tokens, while abstractor only outputs 64 tokens. Note that the computation amount is proportional to the square of the number of tokens.
TABLE XIII: Simply replacing modules (i.e., vision-text connectors and LLMs) has relatively little influence on performance.
Types Architectures Distortion Identification Instant Rating
@@
Vision-text connector Projector 97.9 / 94.7 90.5 / 89.5 82.7 / 81.4 92.7 / 92.4 89.2 / 88.8 96.2 / 95.9 92.1 / 91.9 89.1 / 88.4
Abstractor 97.7 / 94.1 91.3 / 89.3 84.7 / 82.4 93.6 / 93.1 90.5 / 90.0 96.9 / 96.4 92.1 / 91.8 90.0 / 89.6
LLM Vicuna-v1.5-7B 97.9 / 94.7 90.5 / 89.5 82.7 / 81.4 92.7 / 92.4 89.2 / 88.8 96.2 / 95.9 92.1 / 91.9 89.1 / 88.4
Vicuna-v0-7B 96.9 / 93.6 89.8 / 89.3 82.5 / 81.3 92.8 / 92.2 88.2 / 88.1 95.0 / 94.7 91.2 / 91.1 90.5 / 90.2
LLaMA-2-7B 97.0 / 94.0 90.6 / 89.1 81.7 / 81.5 92.6 / 92.0 88.4 / 87.9 94.6 / 94.1 91.8 / 91.1 90.9 / 90.7
Ablation on comparison order. We evaluate the impact of comparison order using the fine-grained dataset of [dataset_compare_order], following its protocol to report consistency / accuracy / correlation. Here, consistency measures prediction stability when the two input images are swapped. As shown in Tab. XIV, our model attains >90% consistency under order reversal, and achieves substantially higher comparison accuracy and score correlation than Q-Instruct and GPT-4V. The Q-Instruct and GPT-4V results are quoted from [dataset_compare_order]. These findings indicate that our model is robust to input-order variations. In addition, we provide the average confidence of our model in Tab. XV. The results show that the confidence of consistent predictions is substantially higher than that of inconsistent ones, highlighting the model’s ability for self-evaluation.
Failure cases. We further examine typical failure cases for a general VLM (i.e., Qwen2.5-VL-7B) and our DepictQA-Wild. Both models sometimes confuse visually similar distortions. In the first row of Fig. 10, they misclassify “compression” (blocky artifacts), “pixelation” (aliasing-induced jaggedness), and “blur” (low-frequency smoothing). In the second row, they confuse “insufficient brightness”, “insufficient saturation”, and “insufficient contrast”, which also produce similar global appearance changes. These results suggest that the fine-grained discrimination ability of current models remains limited.
Figure 10: Failure cases where the model confuses some similar distortions, leading to misclassification.
TABLE XIV: Ablation results of comparison order on fine-grained dataset released in [dataset_compare_order]. We follow [dataset_compare_order] to report the consistency (%) / accuracy (%) / correlation as metrics. The best results are highlighted in bold.
Q-Instruct [qinsight] GPT-4V [gpt4v] DepictQA-Wild
CSIQ (Levels) 11.5 / 8.1 / 0.557 41.9 / 40.2 / 0.906 95.5 / 92.5 / 0.958
CSIQ (Types) 11.7 / 6.9 / 0.416 32.5 / 24.4 / 0.482 90.5 / 69.0 / 0.857
SPAQ (Scores) 44.8 / 23.3 / 0.328 65.3 / 39.8 / 0.448 92.1 / 59.6 / 0.961
TABLE XV: Confidence estimation of consistent and inconsistent comparison results on fine-grained dataset in [dataset_compare_order].
Confidence statistics Consistency Inconsistency
CSIQ (Levels) 0.933 $\pm$ 0.120 0.629 $\pm$ 0.083
CSIQ (Types) 0.860 $\pm$ 0.141 0.611 $\pm$ 0.098
SPAQ (Scores) 0.900 $\pm$ 0.125 0.649 $\pm$ 0.110
V-E Real-world Applications
Assessing web-downloaded images. A practical usage of an IQA model involves assessing the quality of real images. We collect a total of 50 real-world images from the web, featuring diverse contents including animals, plants, faces, buildings, and landscapes. Qualitative results in Fig. 11 and Fig. S8 (in Supp. Mat.) indicate that our method can assess real images with detailed descriptions. More importantly, DepictQA-Wild can describe how the distortions affect the contents. For example, in Fig. 11d, our model first accurately identifies the “severe quantization”, then describes that the quantization “causes banding in the sky and water”, and finally concludes that the quality “is considerably degraded”. We also conduct a user study with 20 participants involved. Participants are instructed to choose the assessment result that is of the highest quality among the test methods. The results are shown in Fig. 7, revealing that our approach stably outperforms baseline methods in aligning human perception.
Figure 11: Qualitative results on assessing web-downloaded images. More results in Fig. S8 of Supp. Mat.
Figure 12: Qualitative results of detailed comparison reasoning on model-processed images.
TABLE XVI: Results on model-processed images. The best results are bold. GPT-4V exhibits variability in ranking, so its results are reported as mean $\pm$ standard deviation.
NIQE ClipIQA MUSIQ ManIQA GPT-4V Ours
[…]
Accuracy $\uparrow$ 45.5 72.7 77.3 66.4 74.5 82.7
Comparison on model-processed images. To develop image restoration models, one often needs to compare the restoration quality of different models. We consider five common image restoration tasks: super-resolution, denoising, JPEG compression artifact removal, motion deblurring, and defocus deblurring. For each task, we collect three to four cutting-edged models in recent years (listed in Tab. XVII), apply them to a correspondingly degraded image, and then manually rank the resultant model-processed images. To find the image considered best by VLMs, we linearly scan the candidates and compare them in pairs. As VLMs’ results are not deterministic and may be sensitive to the presentation order of images, we repeat the linear scan 10 times and randomly shuffle the scan order each time. The average rank of the found best restoration and the accuracy of the paired comparison are reported in Tab. XVI. First, DepictQA-Wild achieves an average rank of 1.20 (1 is the best), outperforming both GPT-4V and score-based methods. Second, though the temperature is set to 0, GPT-4V shows variability with a large standard deviation. Third, model-restored images are generally out-of-distribution for our model, while DepictQA-Wild exhibits excellent generalization ability on these images. For example, the image in Fig. 8 is restored from a noisy image. There is still remnant noise, which is somewhat strange. For such an OOD image, our DepictQA-Wild correctly recognizes it to be inferior, but MANIQA, MUSIQ, and NIQE consider it as the best of the four candidates. Finally, we provide two qualitative results of detailed comparison reasoning on model-processed images in Fig. 12. One compares SwinIR and FBCNN in the JPEG compression artifact removal task, and the other compares MPRNet and MAXIM in the deblur task. Our model can generate a reasonable explanation for the comparison results.
TABLE XVII: Image restoration models used in real-world experiments of comparing model-processed images , including SwinIR [SwinIR], HAT [HAT], X-Restormer [XRestormer], MPRNet [MPRNet], Restormer [Restormer], FBCNN [FBCNN], MAXIM [MAXIM], MPRNet [MPRNet], DRBNet [DRBNet], and IFAN [IFAN]. For FBCNN, “qq=90” means training on the quality factor 90, and “blind” means blind to the quality factor.
[t] Image restoration task Image restoration models Super-resolution SwinIR, HAT, X-Restormer Denoising SwinIR, MPRNet, Restormer, X-Restormer JPEG artifact removal SwinIR, FBCNN (qq=90), FBCNN (blind) Motion deblurring MAXIM, MPRNet, Restormer Defocus deblurring DRBNet, IFAN, Restormer
V-F Complexity and Efficiency
Training cost. The total parameters are 7.11B, including 6.76B for LLM, 0.30B for vision encoder, and 54M for vision abstractor. The trainable parameters are 70M (54M for vision abstractor and 16M for LoRA), constituting only 0.98% of the total parameters. The model is trained on 8 GPUs (RTX A6000). The training is completed in around 22 hours.
Inference cost. The inference latency depends on the response length and it is tested on a single RTX A6000 GPU. For brief tasks task with the short answer prompt (about 2.92 words), the inference time stands at approximately 2.23s / batch=32, transformed to 0.07s / sample. For the assessment reasoning task (75.84 words on average), the inference time is 22.97s / batch=32 (i.e., 0.72s per response). Despite adopting VLMs, our DepictQA-Wild remains deployable on a single consumer GPU (e.g., RTX3090).
VI Conclusions and Limitations
@@
Limitations. First, the fine-grained abilities requiring more high-level perception skills are still unsatisfactory. For example, in Fig. 11c, though identifying noise and pixelation successfully, our model fails to point out that they are respectively located in the left and right parts. One solution is to take the segmentation model to add various distortions to different regions. Second, for the convenience of evaluating, analyzing, and improving the model, we mainly focus on standardized answers. To achieve more flexible responses, LLM rewriting and human annotation can be used to increase linguistic diversity during dataset construction. Third, whether our assessment can be used as feedback to improve the quality of generation or restoration models is still under-explored.
[Uncaptioned image] Zhiyuan You is a Ph.D. candidate from Multimedia Laboratory at The Chinese University of Hong Kong, under the supervision of Prof. Chao Dong and Prof. Tianfan Xue. He obtained both his Bachelor’s and Master’s degrees from Shanghai Jiao Tong University, where he was supervised by Prof. Xinyi Le and Prof. Yu Zheng. His research interests focus on low-level vision based on foundation models.
[Uncaptioned image] Jinjin Gu received his Ph.D. degree from the School of Electrical and Computer Engineering at the University of Sydney in 2024, under the supervision of Prof. Wanli Ouyang and Prof. Luping Zhou. Prior to that, he earned a Bachelor’s degree in Computer Science and Engineering from The Chinese University of Hong Kong, Shenzhen, in 2020. He collaborates closely with Prof. Chao Dong and Prof. Junhua Zhao. His research focuses on computer vision and image processing, with additional interests in the interpretability of deep learning algorithms and the industrial applications of machine learning.
[Uncaptioned image] Xin Cai received the B.S. degree in computer science from University of Chinese Academy of Sciences (UCAS) in 2020 and the M.S. degree from the Institute of Computing Technology(ICT), Chinese Academy of Sciences(CAS),Beijing, China.He is currently pursuing the Ph.D. degree in The Chinese University of Hong Kong (CUHK). His research interests include computer vision, computational photography and machine learning.
[Uncaptioned image] Zheyuan Li is currently a Ph.D. Student at University of Macau. He has also worked as a research intern since 2021 in Shenzhen Institutes of Advanced Technology, Chinese Academy of Sciences. He received his B.E. degree from Northwestern Polytechnical University, Xi’an, in 2022. His research interests include image super-resolution, image restoration and enhancement, and multi-modal low-level-vision model.
[Uncaptioned image] Kaiwen Zhu is currently a Ph.D. student in Shanghai Jiao Tong University supervised by Prof. Chao Dong. He obtained his B.Eng. degree from Shanghai Jiao Tong University in 2024. He has been working as an intern researcher in Shanghai Artificial Intelligence Laboratory since 2023. His research focuses on low-level vision and intelligent agents.
[Uncaptioned image] Chao Dong is a professor at Shenzhen Institutes of Advanced Technology, Chinese Academy of Science (SIAT), and Shanghai Artificial Intelligence Laboratory. In 2014, he first introduced deep learning method – SRCNN into the super-resolution field. This seminal work was chosen as one of the top ten “Most Popular Articles” of TPAMI in 2016. His team has won several championships in international challenges –NTIRE2018, PIRM2018, NTIRE2019, NTIRE2020 AIM2020 and NTIRE2022. He worked in SenseTime from 2016 to 2018, as the team leader of Super-Resolution Group. In 2021, he was chosen as one of the World’s Top 2% Scientists. In 2022, he was recognized as the Al 2000 Most Influential Scholar Honorable Mention in computer vision. His current research interest focuses on low-level vision problems, such as image/video super-resolution, denoising and enhancement.
[Uncaptioned image] Tianfan Xue is a Vice-Chancellor Assistant Professor at the Multimedia Lab in the Department of Information Engineering at the Chinese University of Hong Kong (CUHK). Prior to this, he worked in the Computational Photography Team at Google Research for over five years. He received his Ph.D. degree from the Computer Science and Artificial Intelligence Laboratory (CSAIL) at MIT in 2017. He also holds an M.Phil. degree from CUHK, obtained in 2011, and a Bachelor’s degree from Tsinghua University. His research focuses on computational photography, 3D reconstruction, and generation. His work on bilateral based 3D reconstruction has won SIGGRAPH Honorable mention 2024. He also served as an area chair for WACV, CVPR, NeurIPS, and ACM MM.
-A Dataset Details
-A1 Details of Distortion Library
@@
As stated in Sec. III-B, to facilitate the dataset construction, we design and implement a comprehensive distortion library. Our distortion system contains 12 distortion super-categories in total, with each category consisting of multiple sub-categories. For instance, the “blur” category encompasses “Gaussian blur”, “motion blur”, “lens blur”, etc. In total, there are 35 sub-categories. For each sub-category, there are 5 severity levels: “slight”, “moderate”, “obvious”, “serious”, and “catastrophic”. In this section, we elaborate on our distortion implementations, including the principles, formulas, and severity setup. We also provide one example for each implementation in Fig. S3, with the reference image in Fig. S1.
@@
  • Glass blur. Filter the image using a Gaussian filter, then randomly jitter each pixel in the image by $x$ pixels, and repeat this process $n$ iterations. $\sigma,x,n]\in[(0.7,1,1),(0.9,2,1),(1.2,2,2),(1.4,3,2),(1.6,4,2)]$.
@@
  • Gaussian noise in YCrCb space. Similar to the Gaussian noise in RGB space, this distortion is implemented in YCbCr space, where the value of $(\sigma_{l},\sigma_{r},\sigma_{b})$ can be (0.05, 1, 1), (0.06, 1.45, 1.45), (0.07, 1.9, 1.9), (0.08, 2.35, 2.35), or (0.09, 2.8, 2.8).
@@
Over-sharpen. The reference image $I_{R}$ is firstly processed by a Gaussian blur kernel to generated a blurred image $I_{blur}$. Then the original image is over-sharpened with ${\rm cv2.addWeighted}(I_{R},1+\alpha,I_{blur},-\alpha,0)$, where $\alpha\in[2,2.8,4,6,8]$.
@@
Multi-distortion setups. As discussed in Sec. III-B, multiple distortions may occur simultaneously on the same image in practical usage. First, we observe that humans can identify at most two distortions when three or more are applied, as in Fig. S2, thus we limit the number of applied distortions to two. Second, some distortions could weaken each other’s presentation (e.g., “brighten” weakens “darken”, “blur” weakens “over-sharpen”). Also, certain distortions show similar visual effects (e.g., “pixelate” looks similar to “blur”), making it hard to identify both if applied simultaneously. Hence, to exclude contradictory or similar distortion combinations, we manually review all possible combinations. All feasible distortion combinations are provided in Tab. S1.
TABLE S1: Multi-distortion setting where we show all feasible distortion combinations.
@@
Out-of-distribution setups. In Sec. V-B, we evaluate our model in an out-of-distribution (OOD) setting. Specifically, for a particular category of distortion (e.g., noise), we use some sub-categories (e.g., Poisson noise) during training, and different sub-categories (e.g., impulse noise) for evaluation. Here we provide a detailed split of training distortions and evaluation distortions in Tab. S2.
TABLE S2: Setting of out-of-distribution (OOD) distortion identification.
Category Training distortions Validation distortions
Blur Motion blur, Glass blur, Lens blur, Zoom blur Gaussian blur, Jitter blur
Noise Gaussian noise in YCrCb space, Speckle noise, Spatial correlated noise, Poisson noise Gaussian noise in RGB space, Impulse noise
Compression JPEG compression JPEG2000 compression
Brighten Shift brighten in HSV & RGB spaces, Gamma brighten in HSV space Gamma brighten in RGB space
Darken Shift darken in HSV & RGB spaces, Gamma darken in HSV space Gamma darken in RGB space
Contrast strengthen Contrast strengthen by scaling Contrast strengthen by stretching
Contrast weaken Contrast weaken by scaling Contrast weaken by stretching
Saturate strengthen Saturate strengthen in HSV space Saturate strengthen in YCrCb space
Saturate weaken Saturate weaken in HSV space Saturate weaken in YCrCb space
Quantization Quantization by OTSU method, Quantization by histogram median Quantization by histogram equalization
-A2 Details of Template pool